# ORCID: https://orcid.org/0000-0001-5105-1463

### v2

@@ 3.3 Journal subjects
Table 2 presents a high-level overview of hybrid open access by AJCS subject area using fractionalised counting to account for journals belonging to more than one category. Between 2018 and 2022, most hybrid journals with at least one open access article could be attributed to the Social Sciences category, which also includes Arts and Humanities. However, these journals published the fewest number of articles, whereas Physical Sciences journals recorded the most articles, followed by Health Sciences and Life Sciences. In terms of open access, Physical Sciences journals accounted for more than one-third of the articles published in the five-year period, followed by Health Sciences, Social Sciences and Life Sciences.
@@
Figure 5 presents the relative growth of hybrid open access by subject area between 2018 and 2022. In particular, Social Sciences including Arts and Humanties journals accounted for the strongest growth in the five-years period from 6.4% (n = 8,361) to 23% (n = 51,938), followed by the Life Sciences from 7.6% (n = 15,003) to 18% (n = 39,494) , Health Sciences from 5.3% (n = 18,279) to 16% (n = 63,089) and Physical Sciences from 4.5% (n = 22,364) to 12% (n = 85,428). Growth in Social Sciences category can be largely attributed to transformative agreements. In 2022, two-thirds of open access articles (67%, n = 34,759) were published by the first authors affiliated with participating institutions (see 5B).
Figure 5: Developement of open access in hybrid journals in transformative agreements between 2018 and 2022 by AJCS subject area. The blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. (A) Proportion of open access articles in hybrid journals per year and publisher. (B) Percentage of hybrid open access via agreements per year and publisher. Boxplots (C) show individual open access uptake rates by individual hybrid journals and open access funding per publication year and publisher. The individual outliers are not shown. Note that data on transformative agreements ending before June 2021 were not available for this study.
@@ 3.4 Comparing countries
Between 2018 and 2022, high-income countries almost exclusively dominated hybrid open access publishing through transformative agreements. To discern socio-economic differences, these countries were grouped according to their membership in the Organisation for Economic Co-operation and Development (OECD). During 2018 and 2022, first authors affiliated with institutions from OECD member countries published 602,050 open access articles in hybrid journals, representing 81% of the investigated open access articles. This disparity between OECD nations and other countries becomes even more evident when considering open access through transformative agreements, as 310,712 of 328,957, or 94% of open access articles were associated with such agreements.
Figure 6A shows the development of hybrid open access publishing by country, comparing the OECD area with the BRICS, an intergovernmental organisation, which comprised Brazil, Russia, India, China, and South Africa as of 2022. The residual category “Other” includes the remaining countries. Full counting was applied to account for multiple country affiliations (Hottenrott et al., 2021). From 2018 to 2022, the proportion of open access in hybrid journals increased from 6.1% in 2018 to 26% in 2022. In contrast, BRICS recorded a low uptake, moderately growing from 1.6% in 2018 to 3.7% in 2022.
Despite the rise of open access across OECD countries, the overall publication output decreased sharply, dropping to 786,903 in 2022 after peaking 892,197 articles in 2020. In stark contrast, the number of articles published in hybrid journals by first authors affiliated with institutions from BRICS countries increased steadily over the years, more than doubling from 356,632 in 2018 to 786,903 in 2022. Upon closer examination, this trend can be observed across all of the three largest publishers, although the shift towards BRICS is particularly evident in Elsevier’s hybrid journal portfolio, particularly with regard to articles published in Physical Sciences journals. While OECD publication output in Elsevier’s Physical Sciences journals declined from 112,822 articles in 2018 to 103,766 in 2022, BRICS output increased from 104,654 to 171,713 in the same five-year period. Furthermore, OECD publication output in Health Sciences and Life Sciences journals stagnated after peaking in 2020.
@@
Figure 6: Development of hybrid open access publishing by country. (A) presents the number of articles published in hybrid journals included in transformative agreement, distinguishing between BRICS as of 20222, OECD and other countries. (B) Scatterplot contrasting total articles with open access article volume in 2022, by country. Point size represents the number of articles enabled through transformative agreements. (C) Hybrid open access market share in 2022 by country. In (A) and (C), the blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. Country names are represented as ISO two-letter country codes.
@@
To illustrate the situation in 2022, Figure 6B compares the total publication output with the number of open access articles. With 391,530 articles, China was the most productive country, followed by the United States (n = 268,965) and India (n = 87,428). In contrast, Western and Northern European countries published a considerably high number of open access articles. Particularly, Scandinavian countries, the Netherlands and Austria recorded above-average open access shares, as indicated by the linear trend line. As shown in Figure 6C, transformative agreements contributed to these market positions. Interestingly, the United States had a notable open access market share of 15%, although transformative agreements contributed to a lesser extent. Similarly, China’s open access market share of 7.2% in 2022 was comparable to that of the Netherlands, which was 7.1%.
@@ 4 Discussion
A key finding of this analysis is that transformative agreements maintain market concentration. Specifically, the three largest commercial publishers Elsevier, Springer Nature, and Wiley dominate, particularly with regard to open access provided through transformative agreements. Together, the three publishers accounted for three-fourths of open access articles through transformative agreements, while recording less than half of the total publication volume published between 2018 and 2022. This observation aligns with previous research on shifts in the publishing market following the introduction of funding opportunities for hybrid open access (Butler et al., 2023; Jahn & Tullney, 2016; Shu & Larivière, 2023). Additionally, the results confirm previously observed variations by publisher, with Elsevier exhibiting a different development than Springer Nature and Wiley (Butler et al., 2023; Jahn et al., 2022). However, it must be noted that the focus of transformative agreements on publishers with large hybrid journal portfolios is intentional (Campbell et al., 2022). Because of transformative agreements with a few large publishers, national consortia were able to substantially increase their country’s annual open access article volume (Bosman, Jonge, et al., 2021; Huang et al., 2020; Pinhasi et al., 2021; Taubert et al., 2023; Wenaas, 2022).
@@
Surprisingly, the total publication output of high-income countries belonging to the OECD in hybrid journals declined substantially after peaking in 2020, whereas that of BRICS countries doubled from 2018 to 2022. Because the BRICS expansion can be exclusively attributed to closed access articles, this development has the potential to hinder the transition of academic journal publishing to full open access through transformative agreements and demands discussion. In China, the country with the highest volume of articles in hybrid journals, only limited research funding to pay for open access is available, with expenditures for open access publishing surpassing subscription costs (Shu & Larivière, 2023). Furthermore, the focus of Chinese authors on established journals may contribute to this trend (Zhang et al., 2022). Although China supports Plan S (Schiermeier, 2018), this is not reflected in the data in terms of open access uptake in hybrid journals. As highlighted by Koley & Lala (2023), India faces similar challenges in terms of the availablility of resources to pay for open access. At the same time, access to research literature is a pressing issue, which is addressed by the “Indian one nation, one subscription” policy proposal. However, this policy focuses on centrally negotiated subscriptions and does not entail open access.
@@
This large-scale study provides the first empirical evidence of the influence of transformative agreements on the transition of hybrid journals to full open access. However, several limitations need to be acknowledged. From a data perspective, estimations of open access through transformative agreements were established by linking first author affiliations with publicly available agreement data from cOAlition S, and not through invoicing data, which is usually not shared. Moreover, this study is unable to account for the various types of transformative agreements, particularly with regard to article types and caps that limit the number of open access articles covered. Furthermore, assessing the quality of the OpenAlex and Crossref data used, particularly in terms of affiliations and article types, by combining them with established bibliometric databases such as Scopus and Web of Science was beyond the scope of this analysis (Visser et al., 2021). Overall, the methodology is designed to underestimate, rather than overestimate, the adoption of open access through transformative agreements. In addition, the data used in this study are openly available, along with the code used for this analysis.
@@
This study has practical implications for research funding and libraries. Of primary concern should be the observed differences across countries, particularly the relationship between socio-economic development and open access adoption. From a data perspective, the reporting of open access funding, including transformative agreements, is not harmonised, but often crowd-sourced from various sources. To improve the assessment of transformative agreements, libraries and publishers need to collaborate on standards and services to publicly share information about respective journal portfolios, participating institutions and open access invoicing.

### v3

@@ 3.3 Journal subjects
Table 2 presents a high-level overview of hybrid open access by ASJC subject area using fractional counting to account for journals belonging to more than one category. Between 2018 and 2022, most hybrid journals with at least one open access article could be attributed to the Social Sciences category, which also includes Arts and Humanities. However, these journals published the fewest number of articles, whereas Physical Sciences journals recorded the most articles, followed by Health Sciences and Life Sciences. In terms of open access, Physical Sciences journals accounted for more than one-third of the articles published in the five-year period, followed by Health Sciences, Social Sciences and Life Sciences.
@@
Figure 5 presents the relative growth of hybrid open access by subject area between 2018 and 2022. In particular, Social Sciences including Arts and Humanities journals accounted for the strongest growth in the five-years period from 6.4% (n = 8,361) to 23% (n = 51,938), followed by the Life Sciences from 7.6% (n = 15,003) to 18% (n = 39,494) , Health Sciences from 5.3% (n = 18,279) to 16% (n = 63,089) and Physical Sciences from 4.5% (n = 22,364) to 12% (n = 85,428). Growth in Social Sciences category can be largely attributed to transformative agreements. In 2022, two-thirds of open access articles (67%, n = 34,759) were published by the first authors affiliated with participating institutions (see 5B).
Figure 5: Development of open access in hybrid journals in transformative agreements between 2018 and 2022 by ASJC subject area. The blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. (A) Proportion of open access articles in hybrid journals per year and subject area. (B) Percentage of hybrid open access via agreements per year and subject area. (C) Boxplots show individual open access uptake rates by individual hybrid journals and open access funding per publication year and subject area. The individual outliers are not shown. Note that data on transformative agreements ending before June 2021 were not available for this study.
@@ 3.4 Comparing countries
Between 2018 and 2022, high-income countries almost exclusively dominated hybrid open access publishing through transformative agreements. To discern socio-economic differences, these countries were grouped according to their membership in the Organisation for Economic Co-operation and Development (OECD). Overall, first authors affiliated with institutions from OECD member countries published 602,050 open access articles in hybrid journals, representing 81% of the investigated open access articles. This disparity between OECD nations and other countries becomes even more evident when considering open access through transformative agreements, as 310,712 of 328,957, or 94% of open access articles were associated with such agreements.
@@
Figure 6: Development of hybrid open access publishing by country. (A) presents the relative number of articles published in hybrid journals included in transformative agreement by year, distinguishing between BRICS as of 2022, OECD and other countries. (B) Scatterplot contrasts total articles with open access article volume in 2022, by country and its OECD (purple-colored points) or BRICS (green) membership. Trend line obtained from linear regression, grey area show pointwise symmetric 95% confidence bands. (C) Hybrid open access market share in 2022 by country. In (A) and (C), the blue areas represent open access through transformative agreements, the grey areas depict open access articles where no link to an agreement could be established. The remainder shows closed access articles. Country names are represented as ISO two-letter country codes.
Figure 6A shows the development of hybrid open access publishing by country, comparing the OECD area with the BRICS, an intergovernmental organisation, which comprised Brazil, Russia, India, China, and South Africa as of 2022. The residual category “Other” includes the remaining countries. Full counting was applied to account for multiple country affiliations (Hottenrott et al., 2021). Open access in hybrid journals increased in particular because of publications from first authors affiliated with institutions from the OECD area, from 52,154 in 2018 to 202,787 in 2022. This growth was largely driven by transformative agreements. Their share rose from 5.5% in 2018 (n = 2,859) to 68% (n = 137,815) in 2022. In contrast, BRICS recorded a low uptake, moderately growing from 1.6% in 2018 to 3.7% in 2022.
Despite the rise of open access across OECD countries, the overall publication output decreased sharply, dropping to 786,903 in 2022 after peaking 892,197 articles in 2020, which saw a massive growth in literature related to COVID-19, driven largely by researchers in the United States surpassing China and other countries (Ioannidis et al., 2021). In stark contrast, the number of articles published in hybrid journals by first authors affiliated with institutions from BRICS countries increased steadily over the years, from 356,632 in 2018 to 535,828 in 2022. This resulted in an increase in market share of the BRICS area from 23% to 31% between 2018 and 2022, whereas that of the OECD area decreased from 55% to 46% during the same period. Upon closer examination, this trend can be observed across all of the three largest publishers, although the shift towards BRICS is particularly evident in Elsevier’s hybrid journal portfolio, particularly with regard to articles published in Physical Sciences journals. While OECD publication output in Elsevier’s Physical Sciences journals declined from 112,822 articles in 2018 to 103,766 in 2022, BRICS output increased from 104,654 to 171,713 in the same five-year period. Furthermore, OECD publication output in Health Sciences and Life Sciences journals stagnated after peaking in 2020, which saw an increase in publications due to the impact of the global COVID-19 pandemic, in particular from the United States (Ioannidis et al., 2021).
@@
To illustrate the situation in 2022, Figure 6B compares the total publication output with the number of open access articles. With 391,530 articles, China was the most productive country, followed by the United States (n = 268,965) and India (n = 87,428). In contrast, Western and Northern European countries published a considerably high number of open access articles. Particularly, Nordic countries, the Netherlands and Austria recorded above-average open access shares, as indicated by the linear trend line. As shown in Figure 6C, transformative agreements contributed to these market positions. Interestingly, the United States had a notable open access market share of 15%, although transformative agreements contributed to a lesser extent. Similarly, China’s open access market share of 7.2% in 2022 was comparable to that of the Netherlands, which was 7.1%.
@@ 4 Discussion
A key finding of this analysis is that transformative agreements maintain market concentration. Specifically, the three largest commercial publishers Elsevier, Springer Nature, and Wiley dominate, particularly with regard to open access provided through transformative agreements. Together, the three publishers accounted for three-fourths of open access articles through transformative agreements, while recording about half of active hybrid journals included in transformative agreements between 2018 and 2022. This observation aligns with previous research on shifts in the publishing market following the introduction of funding opportunities for hybrid open access (Butler et al., 2023; Jahn & Tullney, 2016; Shu & Larivière, 2023). Additionally, the results confirm previously observed variations by publisher, with Elsevier exhibiting a different development than Springer Nature and Wiley (Butler et al., 2023; Jahn et al., 2022). However, it must be noted that the focus of transformative agreements on publishers with large hybrid journal portfolios is intentional (Campbell et al., 2022). Because of transformative agreements with a few large publishers, national consortia were able to substantially increase their country’s annual open access article volume (Bosman, Jonge, et al., 2021; Brayman et al., 2024; Huang et al., 2020; Pinhasi et al., 2021; Taubert et al., 2023; Wenaas, 2022; Widding, 2024).
@@
Surprisingly, the total publication output of high-income countries belonging to the OECD in hybrid journals declined substantially after peaking in 2020, whereas that of BRICS countries continuously grew from 2018 to 2022. Because the BRICS expansion can be exclusively attributed to closed access articles, this development has the potential to hinder the transition of academic journal publishing to full open access through transformative agreements and demands discussion. In China, the country with the highest volume of articles in hybrid journals, only limited research funding to pay for open access is available, with expenditures for open access publishing surpassing subscription costs at some universities (Shu & Larivière, 2023). Furthermore, the focus of Chinese authors on established journals may contribute to this trend (Zhang et al., 2022). Although China supports Plan S (Schiermeier, 2018), this is not reflected in the data in terms of open access uptake in hybrid journals. Market observers also do not expect a broad implementation of transformative agreements in China in near future (Owens, 2024). As highlighted by Koley & Lala (2023), India faces similar challenges in terms of the availability of resources to pay for open access. At the same time, access to research literature is a pressing issue, which is addressed by the “Indian one nation, one subscription” policy proposal. However, this policy focuses on centrally negotiated subscriptions and does not entail open access.
@@
This large-scale study provides the first empirical evidence of the influence of transformative agreements on the transition of hybrid journals to full open access. However, several limitations need to be acknowledged. From a data perspective, estimations of open access through transformative agreements were established by linking author affiliations with publicly available agreement data from cOAlition S, and not through invoicing data, which is usually not shared. Because data on corresponding authors, who are typically responsible for facilitating open access publication, were not fully supported by OpenAlex at the time of this study, first author affiliations were used instead. This study is also unable to account for the various types of transformative agreements due to a lack of data, especially regarding article types and caps that limit the number of open access articles covered. Furthermore, assessing the quality of the OpenAlex and Crossref data used, particularly in terms of affiliations, corresponding authorship, and article types, by combining them with established bibliometric databases such as Scopus and Web of Science was beyond the scope of this analysis (Chinchilla-Rodríguez et al., 2024; Visser et al., 2021). It must therefore be emphasised that, unlike evaluations from national consortia that could make use of invoices (Brayman et al., 2024), this global overview can only provide an estimate of the support for open access through transformative agreements. Despite these limitations, the methodology was designed to underestimate, rather than overestimate, the adoption of open access through transformative agreements. To promote transparency, the data used in this study, along with the code used for this analysis, are openly available.
@@
This study has practical implications for research funding and libraries. One concern should be the observed differences across countries, particularly the relationship between socio-economic development and open access adoption. Between 2018 and 2022, there was a notable increase in the number of closed access articles published in hybrid journals by authors from countries such as China and India, where transformative agreements and open access funding options were not widely available. In contrast, the introduction of country-wide transformative agreements in numerous OECD countries led to a substantial increase in open access, while their share of total articles in hybrid journals decreased. This imbalance, whereby open access uptake depends largely on a few countries, makes it less likely that collective action towards a large-scale transition of hybrid journals to full open access through transformation agreements will succeed within the next years. Rather, and in accordance with recent British and Scandinavian policy recommendations (Brayman et al., 2024; Widding, 2024), it emphasises the importance of ongoing evaluation and adaptation of open access strategies in view of global publishing trends and related challenges to equity and journal quality (Rasmussen, 2023; Ross-Hellauer et al., 2022).
From a data perspective, research institutions and funders need to push towards price transparency in scholarly publishing to enable these evaluations. The reporting of open access funding, including transformative agreements, is also not harmonised, but often crowd-sourced from various sources. To improve the assessment of transformative agreements, libraries and publishers should collaborate on standards and services to publicly share information about respective journal portfolios, participating institutions and open access invoicing, for example in the international context of ESAC and the Barcelona Declaration on Open Research Information. Such collaboration would help overcome the limitations of current approaches that derive funding for open access from authorship data.